Susan Marciano invested part of her $23,000 bonus in a fund that paid 10% profit and invested the rest in stock that suffered a 3% loss. find the amount of each investment if her overall net profit was $2,040.
what's the amount invested at10%? whats the amount invested in stock is?

Answer by richwmiller(17219)

Total amount of money invested: $23000
x+y=23000,
Total yearly interest for the two accounts is: $2040
0.1*x+-0.03*y=2040
x=23000-y
Substitute for x
0.1*(23000-y)+-0.03*y=2040
Multiply out
2300-0.1y+-0.03*y=2040
Combine like terms.
-0.13*y=-260
Isolate y
y=$ 2000.00 at -3%
x=23000-y
Calculate x
x=$21000.00 at 10%
Check
0.1*21000+-0.03*2000=2040
2100+-60=2040
2040=2040
If this statement is TRUE and neither x nor y is negative then all is well